# Boosting Memory Efficiency in Transfer Learning for High-Resolution Medical Image Classification

### v2

@@ IV-H Visualization
To verify the important tokens selection and the mechanism of the fusion modules, we visualize the selected tokens of the top layer and the attention scores on them in the corresponding fusion module across four modalities. As shown in Fig. 9, the selected important tokens are highlighted in green boxes. It can be observed that the selected tokens cover most discriminative regions, such as lesions and the optic disc in fundus images. Please note that ViT employs a multi-head mechanism, and the visualization is only for one of the multiple heads. This multi-head mechanism ensures that the important tokens selection can convey sufficient diagnostic clues to the side network. Additionally, we visualize the attention that the fine-grained prompts pay to the selected tokens. It can be observed that the fine-grained prompts further refine the selection, reducing redundancy and summarizing them into a small set of prompts for the side network. This refinement process ensures that only the most relevant information is conveyed, enhancing the efficiency and effectiveness of the transfer learning process.
@@
Figure 9: Visualization of the selected important tokens and attention scores from the fine-grained prompts.
@@ V Discussion
With the success of large-scale pre-trained models, transfer learning has become a key technique to apply pre-trained knowledge in real-world applications. PETL has been proposed to efficiently adapt LPMs to new downstream tasks. However, existing PETL methods mainly focus on reducing the number of learnable parameters, and the improvements in memory efficiency are not as effective as those in parameter efficiency. Medical image analysis is a domain that often requires high-resolution inputs to ensure that tiny diagnostic clues are clearly exposed. However, increasing the input resolution dramatically increases memory consumption during training. As shown in Fig. 2, even the most memory-efficient PETL method, linear probing, is not feasible when the input resolution is $1280\times 1280$ on a GPU with a 48GB memory size. Our proposed FPT+ significantly improves memory efficiency, making training at $2048\times 2048$ possible on a single GPU, with the potential for even higher resolutions based on the trend shown in the figure.
Although FPT+ makes the training of super high-resolution images feasible on one GPU, we notice that the performance saturates or even slightly drops when the input resolution exceeds $1024\times 1024$. There may be two plausible reasons: (1) The LPM used is pre-trained at a standard resolution of $224\times 224$, and its feature extraction ability degrades when the resolution is significantly different from the pre-trained resolution; (2) The datasets we use may have already been resized and compressed by their owners, allowing users to download them from the website. Hence, increasing the input size to go beyond the original image size shall not provide any additional information.
High-resolution image classification is not the only field that benefits from this memory efficiency. FPT+ can also be applied to ViT variants with smaller patch sizes. ViTs with smaller patch sizes generally have more powerful abilities to extract fine-grained semantics. However, smaller patch sizes result in longer input sequences for the same input resolution. FPT+ may be a good solution to attain the best balance. The memory efficiency of FPT+ also allows for larger batch sizes during training, which can accelerate the training process and potentially improve the stability and performance of transfer learning. Additionally, there are settings in natural image analysis that also require high resolution, such as satellite imagery [1]. The generalizability of FPT+ to natural images remains unexplored, which is one of our future works.
@@ VI Conclusion
In this paper, we extend our previous work FPT [23] to FPT+. FPT+ significantly improves the memory efficiency over existing PETL methods, particularly in the high-resolution context commonly encountered in medical image analysis. To reduce memory consumption, we first enhance the design of side tuning with asymmetric input. To effectively adapt pre-trained knowledge from the LPM, we introduce fine-grained prompts and fusion modules to summarize and process intermediate features of the LPM and effectively convey them to the side network. To further reduce memory utilization, we propose the selection of important tokens and the preloading of pre-trained features. By integrating all these components, we present a parameter and memory-efficient transfer learning method for high-resolution medical image classification. The effectiveness of FPT+ is successfully established across eight medical image datasets of diverse sizes, modalities, and complexities. Experimental results show that FPT+ outperforms all other SOTA PETL methods while maintaining the best parameter and memory efficiency.

### v3

@@ IV-J Visualization
To verify the important tokens selection and the mechanism of the fusion modules, we visualize the selected tokens of the top layer and the attention scores on them in the corresponding fusion module across four modalities. As shown in Fig. 10, the selected important tokens are highlighted in green boxes and clinically relevant areas are highlighted by red bounding boxes. It can be observed that the selected tokens cover most discriminative regions, such as lesions and the optic disc in fundus images. Please note that ViT employs a multi-head mechanism, and the visualization is only for one of the multiple heads. This multi-head mechanism ensures that the important tokens selection can convey sufficient diagnostic clues to the side network. Additionally, we visualize the attention that the fine-grained prompts pay to the selected tokens. It can be observed that the fine-grained prompts further refine the selection, reducing redundancy and summarizing them into a small set of prompts for the side network. This refinement process ensures that only the most relevant information is conveyed, enhancing the efficiency and effectiveness of the transfer learning process.
@@
Figure 10: Visualization of the selected important tokens and attention scores from the fine-grained prompts for representative cases. Clinically relevant areas are highlighted by red bounding boxes.
@@ V Discussion
With the success of large-scale pre-trained models, transfer learning has become a key technique to apply pre-trained knowledge in real-world applications. PETL has been proposed to efficiently adapt LPMs to new downstream tasks. However, existing PETL methods mainly focus on reducing the number of learnable parameters, and the improvements in memory efficiency are not as effective as those in parameter efficiency. Medical image analysis is a domain that often requires high-resolution inputs to ensure that tiny diagnostic clues are clearly exposed. However, increasing the input resolution dramatically increases memory consumption during training. As shown in Fig. 2, even the most memory-efficient PETL method, linear probing, is not feasible when the input resolution is $1280\times 1280$ on a GPU with a 48GB memory size. Our proposed FPT+ significantly improves memory efficiency, making training at $048\times 2,$ 048 × 2 , 048 possible on a single GPU, with the potential for even higher resolutions based on the trend shown in the figure.
Although FPT+ makes the training of super high-resolution images feasible on one GPU, we notice that the performance saturates or even slightly drops when the input resolution exceeds $024\times 1,$ 024 × 1 , 024. There may be two plausible reasons: (1) The LPM used is pre-trained at a standard resolution of $224\times 224$, and its feature extraction ability degrades when the resolution is significantly different from the pre-trained resolution; (2) The datasets we use may have already been resized and compressed by their owners, allowing users to download them from the website. Hence, increasing the input size to go beyond the original image size shall not provide any additional information.
High-resolution image classification is not the only field that benefits from this memory efficiency. The memory efficiency of FPT+ allows for larger batch sizes during training, which can accelerate the training process and potentially improve the stability and performance of transfer learning. Additionally, there are settings in natural image analysis that also require high resolution, such as satellite imagery [1]. The generalizability of FPT+ to natural images remains unexplored, which is one of our future works.
@@ VI Conclusion
In this paper, we extend our previous work FPT to FPT+. FPT+ significantly improves the memory efficiency over existing PETL methods, particularly in the high-resolution context commonly encountered in medical image analysis. To reduce memory consumption, we first enhance the design of side tuning with asymmetric input. To effectively adapt pre-trained knowledge from the LPM, we introduce fine-grained prompts and fusion modules to summarize and process intermediate features of the LPM and effectively convey them to the side network. To further reduce memory utilization, we propose the selection of important tokens and the preloading of pre-trained features. By integrating all these components, we present a parameter and memory-efficient transfer learning method for high-resolution medical image classification. The effectiveness of FPT+ is successfully established across eight medical image datasets of diverse sizes, modalities, and complexities. Experimental results show that FPT+ outperforms all other SOTA PETL methods while maintaining the best parameter and memory efficiency.